New furniture has a cost of $34000 and a 7 year depreciation life. The estimated Salvage Valus is Zero at the end of 7 years. Determine the annual depreciation amounts using the 200% Declining Balance with Switchover to Straight line. *Tabulate the annual depreciation amounts and the book value at the end of each year
Answer #1

Sol : furniture cost =$34000 depreciation life =7 year

Salvage value =0

Under the straight line depreciation method

Annual depreciation =34000/7=4857.14 (0 salvage value)

  • Here 4857.14 is depreciation expense. Now we have to find that what percent is it of total amount ($3 4000)

​​​ = 4857.14/34000 =0.1429 =14.29%

  • Now for 200% method, multiply (2×14.29)=28.58

Depreciation calculation using the 200 percent double declining balance method :-

Year applicable %dep. Rate Start carrying value

Depreciation expence

  

Ending carrying value
1 28.58 34000 9717.2 24282.8
2 28.58 24282.8 6940.02 17342.78
3 28.58 17342.78 4956.57 1238.21
4 28.58 12386.21 3539.98 8846.23
5 28.58 8846.23 2528.25 6317.98
6 28.58 6317.98 1805.67 4512.3

7 28.58 4512.3 1289.62 0

+3222.68

Special adjustment for final year.

The final year adjustment was calculated because the carrying value at the end of 7 year is zero.